# Rotate, Clip, and Partition: Towards W2A4KV4 Quantization by Integrating Rotation and Learnable Non-uniform Quantizer

### v1

@@ Abstract
We propose Rotate, Clip, and Partition (RCP), a Quantization-Aware Training (QAT) approach that first realizes extreme compression of LLMs with W2A4KV4(2-bit weight, 4-bit activation, and 4-bit KV-cache) configuration. RCP integrates recent rotation techniques with a novel non-uniform weight quantizer design, by quantitatively analyzing the impact of random rotation on 2-bit weight quantization. Our weight quantizer features Learnable Direct Partitioning (LDP), which introduces learnable parameters to directly learn non-uniform intervals jointly with LLM weights. We also present a specialized GPU kernel that supports GEMV on non-uniform W2A4. Experiments show that RCP can compress LLaMA-2-7B to W2A4KV4 with a loss of only 2.84 WikiText2 ppl and 5.29 times reduced memory footprint. Furthermore, RCP can quantize challenging mobile-targeted LLaMA-3.2 models and domain-specific WizardCoder-7B and MetaMath-7B with no critical problems such as convergence failure and repetition. Code will be made available at blind_review.
@@
Euntae Choithanks: these authors contributed equally., Sumin Song11footnotemark: 1, Woosang Lim, Sungjoo Yoo Seoul National University euntae.choi175@gmail.com, songsm921@snu.ac.kr, ftyg656512@snu.ac.kr, sungjoo.yoo@gmail.com
@@ 1 Introduction
Large language models (LLMs) have made significant advancements, but their growing size and resource demands create challenges for deployment across data centers and mobile devices. To address these constraints, extensive research efforts have focused on improving the efficiency of LLM serving through various strategies. Quantization, one of the various methods, has emerged as a particularly straightforward and effective method for reducing memory consumption and inference costs without severely compromising accuracy. By lowering the numerical precision of parameters and intermediate representations, quantization leverages hardware capabilities such as half-precision and integer tensor cores.
[…]
@@
To address these issues, rotation techniques Ashkboos et al. (2024b); Liu et al. (2024b) have proven effective at W4A4KV4. Nevertheless, extending these methods to even lower bit widths, such as W2 or W3, remains insufficient, as the severe information bottleneck cannot be overcome with a uniform quantization and PTQ, ultimately resulting in a loss of generation quality.
Figure 2: A chain of thought (CoT) reasoning example from GSM8K Cobbe et al. (2021) benchmark, conducted by two differently quantized MetaMath-7B Yu et al. (2023) models. The result on the left is from a state-of-the-art QAT method BitDistiller Du et al. (2024). On the right, our proposed RCP is applied. Both methods employ exactly the same 4-bit quantization setting for activation and KV-cache.
In this work, we propose a Quantization-Aware Training (QAT) approach called Rotate, Clip, and Partition (RCP), which integrates rotation techniques with non-uniform quantization strategies to achieve extreme compression of LLMs, specifically realizing W2A4KV4 configurations. Our method jointly optimizes quantization parameters and model weights by systematically incorporating rotation and non-uniform group quantization. We recover meaningful model performance and demonstrate that even 2-bit weight precision, once deemed unattainable, is now within reach. We further extend our techniques to more challenging models, such as LLaMA-3.2 AI@Meta (2024), as well as smaller, mobile-targeted architectures, thereby broadening the applicability of these methods. To facilitate practical deployment, we design efficient W2A4 inference kernels for decoding phase on modern GPUs, thus enabling real-time, resource-efficient LLM serving at extreme compression levels.
@@
  • We propose a QAT algorithm that successfully leverages the benefit of random rotation to first achieve W2A4KV4 quantization of LLMs.
  • We demonstrate the scalability of our approach to more challenging (e.g., LLaMA-3.2) and smaller, mobile-targeted models.
  • Additionally, we develop efficient W2A4 inference kernels for the decoding stage on modern GPUs, enabling real-time, resource-efficient LLM serving under extremely low-bit precisions.
@@ 2.1 Random Rotation for LLM Quantization
As shown in Fig. 4, $R_{1}$ rotates all transformer decoder layers’ input and output activations and its inverse ($R_{1}^{T}$) is usually fused into the neighboring model weights. $R_{2}$ and $R_{4}$ both require online rotation during inference as they target the intermediate activations in the MHA and FFN layers, respectively. $R_{2}$ is decomposed into two orthogonal matrices: one with head dimension size ($R_{H}$), applied during the V projection, and another with the number of heads ($R_{H}^{\prime}$), applied to the self-attention activation. Its transpose ($R_{2}^{T}$) is applied as a whole to the out-projection in the self-attention layers. $R_{3}$ rotates query (Q) and key (K) vectors after RoPE, ensuring that the KV-cache can be compressed without affecting the self-attention output.
2.2 Asymmetric Quantization and Clipping
Asymmetric quantization uses min/max of data to match data’s asymmetric distribution. This approach can offer higher quantization accuracy than symmetric quantization. The asymmetric quantization equation can be formulated as follows:
@@ 2.2 Asymmetric Quantization and Clipping
$\begin{split}\mathbf{W_{q}}=\mathrm{clamp}(\lfloor\frac{\mathbf{W}}{h}\rceil+z,0,2^{N}-1),\\ \mathrm{where}\,\,h=\frac{\max(\mathbf{W})-\min(\mathbf{W})}{2^{N}-1},\\ z=-\lfloor\frac{\min(\mathbf{W})}{h}\rceil\end{split}$ (1)
Clipping is an optimization technique that truncates outlier values to a predetermined range before quantization. It improves quantization precision within the compressed range, preserving performance while increasing compression ratios. Our proposed method, RCP is based on the learnable weight clipping (LWC) Shao et al. (2024), an asymmetric quantization to obtain asymmetric clipping values that minimize quantization error for each linear layer and then uses them as initial values for learnable clipping $f$ as follows:
$f(\textbf{W},\beta,\gamma)=clip(\textbf{W},\sigma(\beta)\text{min}(\textbf{W}),\sigma(\gamma)\text{max}(\textbf{W}))$ (2)
where $\beta$ and $\gamma$ are learnable parameters for clipping and $\sigma$ is the sigmoid function.
3 Observation: Rotation Sharpens Weight Distribution
(a) Per-group kurtosis measured on the 15th key projection weight in LLaMA-2 7B model before and after random Hadamard transform with layernorm fusion.
(b) Per-group kurtosis measured on the input activation of the 15th key projection weight in the same manner as 3(a) in LLaMA-2 7B model. Due to a large difference of average value, the y axis is presented in log scale. Since almost all groups had small negative kurtosis value after transform, we take the absolute value of them.
Figure 3: Kurtosis analysis results to demonstrate the effect of random rotation on model weight and activation.
As shown in Fig. 1, simply applying rotation techniques before LLM quantization can’t guarantee the performance in extremely low-bit scenarios like W2A4. Building on SmoothQuant Xiao et al. (2023)’s insights, outlier suppression methods improve quantization possibility in the activation space while accepting modest weight space compromises. But recent rotation based PTQ techniques predominantly emphasize the advantages in the activation space, not the disadvantages in the weight space. While rotation-based PTQ techniques work well for moderate quantization up to W4, at W2, the quantization possibility significantly drops.
In this section, we analyze quantization possibility after random Hadamard transformation in both weight and activation spaces in LLaMA-2 7B model. Kurtosis, which measures the fourth standardized moment of a distribution, serves as a key indicator for evaluating outlier frequency and concentration Lee et al. (2024) that intensify the difficulties in quantization. We selected the 15th key projection module for analysis because it showed one of the highest mean squared errors (MSE) when comparing the distributions before and after transformation among all modules. We analyze kurtosis of weight and input activation in group size of 128 following their channels to fully consider the effects of groupwise quantization configuration used in most quantization methods.
Our analysis in Fig. 3(b) reveals that rotation effectively distributes activation values along their channel direction, similar to methods like SliceGPT Ashkboos et al. (2024a), QuaRot Ashkboos et al. (2024b) and PrefixQuant Chen et al. (2024a). The original space shows infinite kurtosis due to extreme outliers, while the rotated space shows predominantly near-zero values. From the perspective of per-channel uniform quantization in the activation, the transformation reduces quantization error compared to the original space.
Conversely, our analysis in Fig. 3(a) shows contrasting circumstances. After going through essential processes such as LayerNorm Fusion and Random Hadamard Transformation, an initially platykurtic distribution with mean kurtosis near 1 becomes more leptokurtic with mean kurtosis near 3.75. These observations indicate that the weight distribution deviates further from the uniform distribution, leading to significant quantization errors when using W2 uniform quantizers, as most weights cluster in two center quantization bins. This finding motivates the development of non-uniform quantizers to optimize bin width distributions. This analysis also explains why FP3 variants outperform INT3 quantization in W3 regime, as NF3’s approach better handles bell-shaped weight distributions, as demonstrated in AFPQ Zhang et al. (2023).
@@ 4 Methodology
We propose a non-linear weight quantization (Section 4.1), an efficient dequantization for non-linear weight partitions (Section 4.2), and a GPU inference kernel for extreme low-bit weights (Section 4.4).
Figure 4: An overview of our QAT-based KD framework of RCP. RCP performs rotation-aware clipping, followed by learnable non-uniform quantization (LDP). During this process, knowledge distillation is employed as the training method.
4.1 Differentiable Non-uniform INT2 Group Quantizer
Rotation-aware Clipping Initialization
We initialize clipping parameters in a rotation-aware manner as follows:
$\operatorname*{argmin}_{\beta,\gamma}\|Q(\textbf{W}_{R})\textbf{X}_{R}-\textbf{W}_{R}\textbf{X}_{R}\|^{2}\vspace{-0.2cm}$ (3)
where $\textbf{W}_{R}=R^{T}_{front}f(\textbf{W},\beta,\gamma)R_{rear}$ is the weight matrix obtained by applying appropriate rotations after the clipping $f$ in Eqn. 2. For instance, $R_{front}$ and $R_{rear}$ are set to $R_{1}$ and $I$ for up and gate projections, respectively. Rotation configuration for the other types of weights can be found on the left side of Fig. 4.
Learnable Direct Partitioning
@@ 4.1 Differentiable Non-uniform INT2 Group Quantizer
Figure 5: A diagram of Learnable Direct Partitioning (LDP).
The key component of our proposed RCP is a weight quantizer module named learnable direct partitioning (LDP). In order to realize asymmetric weight quantization, we determine value range by LWC and the partitions within the value range by LDP. LDP starts by normalizing the weight W by a dynamic range $h=\sigma(\gamma)\text{max}(\textbf{W})-\sigma(\beta)\text{min}(\textbf{W})$ determined by the LWC. In LDP, two partitioning variables $s_{1}$ and $s_{2}$ are introduced per weight group [^2] For brevity, we do not use a separate notation for group quantization unless necessary since our method can be applied likewise simply by reshaping a weight so that the last dimension becomes the group size., which split the range $h$ into three subsections.
$\begin{split}p_{1}=\sigma(s_{1}),p_{2}=(1-p_{1})\sigma(s_{2}),\\ p_{3}=(1-p_{1})(1-p_{2})\end{split}$ (4)
As shown in Fig. 5 and Eqn. 4, $s_{1}$ defines the portion of the first partition $p_{1}$ in the whole dynamic range $h$. $s_{2}$ defines the portion of the second partition $p_{2}$ from the remaining range $(1-p_{1})h$. The third one is trivially calculated. This formulation guarantees that i) the dynamic range $h$ is seamlessly filled out, ii) each portion is constrained between 0 and 100% via the sigmoid re-parametrization, and iii) no matter how the parameters are updated, the order of the partitions stays the same. The partitioning parameters are initialized as $s_{1}=\sigma^{-1}(1/3)$ and $s_{2}=\sigma^{-1}(1/2)=0$ to evenly split the quantization range in the beginning. The AWQ-styled grid search Lin et al. (2024a) can also be used to find the optimal partition widths; however, the computational burden will grow exponentially since we have to iterate over a four-dimensional loop (two for LWC and two for LDP).
The quantization process of our LDP can then be derived as follows:
$\textbf{W}_{q}=\text{u}(\frac{\textbf{W}}{h}-t_{1})+\text{u}(\frac{\textbf{W}}{h}-t_{2})+\text{u}(\frac{\textbf{W}}{h}-t_{3})$ (5)

### v2

@@ Abstract
We propose Rotate, Clip, and Partition (RCP), a Quantization-Aware Training (QAT) approach that first realizes extreme compression of LLMs with W2A4KV4 (2-bit weight, 4-bit activation, and 4-bit KV-cache) configuration. RCP integrates recent rotation techniques with a novel non-uniform weight quantizer design by theoretically and empirically analyzing the impact of rotation on the non-uniformity of weight distribution. Our weight quantizer, Learnable Direct Partitioning (LDP), introduces learnable parameters to directly learn non-uniform intervals jointly with LLM weights. We also present a GPU kernel supporting GEMV on non-uniform W2A4 as proof of concept. Experiments show that RCP can compress LLaMA-2-7B to W2A4KV4 with a loss of only 2.84 WikiText2 PPL and 5.29 times reduced memory footprint. Furthermore, RCP can quantize challenging mobile-targeted LLaMA-3.2 models and domain-specific WizardCoder-7B and MetaMath-7B with no critical problems such as convergence failure and repetition. Code is available at https://github.com/songsm921/RCP.
@@
Euntae Choithanks: indicates equal contribution, Sumin Song11footnotemark: 1, Woosang Lim, Sungjoo Yoothanks: indicates corresponding author Seoul National University euntae.choi175@gmail.com, songsm921@snu.ac.kr, ftyg656512@snu.ac.kr, sungjoo.yoo@gmail.com
@@ 1 Introduction
Large language models (LLMs) have made significant advancements, but their growing size and resource demands create challenges for deployment across data centers and mobile devices. To address these constraints, extensive research efforts have focused on improving quantization algorithms.
Notably, rotation-based Post-Training Quantization (PTQ) Ashkboos et al. (2024b); Liu et al. (2024b); Lin et al. (2024a) showed remarkable improvement on W4A4KV4111We call $l$-bit weight, $m$-bit activation, and $n$-bit KV-cache WllAmmKVnn like W2A4KV4. quantization, and state-of-the-art Quantization-Aware Training (QAT) Du et al. (2024); Chen et al. (2024) made extreme weight quantization possible via careful design of datasets and training procedures.
In this work, we propose Rotate, Clip, and Partition (RCP), a rotation-based QAT algorithm, to push the boundaries of extremely low-bit compression. Based on empirical and theoretical analysis, we draw our main insight that rotating LLM weights has two effects at once: eliminating outliers and increasing the non-uniformity of the weight distribution. The key component of RCP is Learnable Direct Partitioning (LDP), which is a fully differentiable non-uniform weight quantizer working in three steps: 1) quantization range setup with learnable clipping parameters Shao et al. (2024); 2) non-uniform quantization via splitting the quantization range with learnable partitioning parameters; 3) non-uniform dequantization that maps the quantized weights to real-valued grids.
@@
RCP is the first to enable challenging W2A4KV4 and W3A4KV4 quantization on common LLM models, significantly outperforming QuaRot Ashkboos et al. (2024b) and BitDistiller Du et al. (2024). Especially, as we present in Fig. 1, RCP also works on small and mobile-targeted LLaMA-3.2 models AI@Meta (2024) that are harder to quantize. Since there is no available hardware that supports both LUT inference for non-uniform quantization and specialized acceleration for 4-bit activation, we design an accelerated GEMV kernel in CUDA as a proof of concept. Our kernel can reduce the memory footprint up to 5.29 times with a latency lower than the FP16 PyTorch Paszke et al. (2019) and INT4 QuaRot implementation.
@@
  • We provide empirical and theoretical analysis on how rotation interacts with weight distribution and poses difficulty on extreme W2A4KV4 quantization.
  • To address this issue, we introduce RCP, a quantization algorithm that takes the best from rotation and QAT via LDP, a novel fully learnable non-uniform quantizer.
  • We provide extensive experiments to show RCP achieves state-of-the-art W2A4KV4 and W3A4KV4 quantization for the first time.
@@ 2.1 Random Rotation for LLM Quantization
Figure 2: A diagram of the rotation process in a transformer decoder layer.
@@
[…]
2.2 Asymmetric Integer Quantizer
The commonly used asymmetric integer quantization function is defined in Eqn. 1.
@@ 2.2 Asymmetric Integer Quantizer
$\begin{split}Q(\mathrm{\mathbf{W}})=\mathrm{clip}(\lfloor\frac{\mathbf{W}}{s}\rceil+z,0,2^{N}-1),\\ \mathrm{where}\,\,h=\max(\mathbf{W})-\min(\mathbf{W}),\;s=\frac{h}{2^{N}-1}\\ z=-\lfloor\frac{\min(\mathbf{W})}{h}\rceil,\end{split}$ (1)
where $N$ is the number of bits, $h$ is the quantization range, $s$ is the step size, and $z$ is the zero-point. This general formulation is applicable to various settings, including per-tensor, per-channel, and group-wise quantization, via adapting the computation of $h$, $s$, and $z$.
3 Motivation
In this section, we confirm the difficulty of W2A4KV4 quantization by empirical evaluations and justify our key design of non-uniform weight quantizer (in Section 4.3) through theoretical analysis on the effect of the rotation technique on weight distribution.
Method $R$ Language Reasoning
WikiText2 Coding Math
QuaRot 12772.03 0 0.002
BitDistiller 17.40 3.5 5.39
BitDistiller 8.93 6.09 0.16
RCP 8.31 23.20 40.16
Table 1: Evaluation results on WikiText2, HumanEval, and GSM8K under W2A4KV4. The evaluations are conducted using LLaMA-2 7B for WikiText2 (perplexity), WizardCoder 7B for HumanEval (pass@1), and MetaMath 7B for GSM8K (pass@1). The column $R$ indicates whether rotation is applied.
Existing algorithms can fail on W2A4KV4
As shown in Table 1, we first observe that QuaRot and BitDistiller fail under W2A4KV4, particularly on language modeling and reasoning tasks. This demonstrates their limitations: QuaRot effectively mitigates activation outliers but fails to handle extreme low-bit weight quantization. BitDistiller is able to address weight quantization but remains vulnerable to the activation outliers.
Naturally, we conceptualized combining rotation and QAT approaches and conducted experiments with all rotation matrices applied to the BitDistiller implementation. As indicated in the "BitDistiller w/ Rotation" rows in Table 1, language modeling performance was recovered to some extent; however, reasoning capabilities remained non-deployable.
Figure 3: A two-dimensional histogram comparing the increase in output activation’s quantization error and the increase in the averaged group-wise weight kurtosis. The last down-projection layer of the LLaMA-2-7B model is used. The bright blue horizontal line indicates the average increase of the quantization error. We use mean absolute error $QErr(\mathbf{W},\mathbf{X})=|\mathbf{X}(Q(\mathbf{W})-\mathbf{W})|$ and $\mathbf{W}_{rc}$ follows the definition in Eqn. 6.
Relation between rotation and non-uniformity
To explain why such a naive application of rotation to QAT fails, we first theoretically analyze how rotation affects the excess kurtosis of the weight distribution. The excess kurtosis is the shifted fourth standardized moment defined as follows:
${Kurt}(X)=\frac{\mu_{4}}{\sigma^{4}}-3,$ (2)
where $\mu_{4}$ is the fourth moment and $\sigma$ is the standard deviation. Larger excess kurtosis indicates a distribution 1) contains numerous outliers and/or 2) is more peaked around its center (i.e., more non-uniform, which is hard to quantize).
Our claim is that the Hadamard matrix (used as rotation) increases the excess kurtosis of a platykurtic distribution [^2] A distribution $X$ is platykurtic when $Kurt(X)<0$, which we empirically observed to be true for most of the LLM weights.
Lemma 1.
Let $\mathbf{X}=(X_{1},X_{2},\ldots,X_{n})^{T}$ be a random vector whose components are i.i.d. with finite fourth moment $\mu_{4}$, mean $\mu$, variance $\sigma^{2}$, and negative excess kurtosis ($\text{Kurt}(X_{i})<0$). Let $\mathbf{H}_{n}$ denote the normalized $n\times n$ Hadamard matrix with elements $\pm\frac{1}{\sqrt{n}}$. Then, for the transformed vector $\mathbf{Y}=\mathbf{H}_{n}\mathbf{X}$, the following holds:
$\text{Kurt}(Y_{i})>\text{Kurt}(X_{i})$
for all $i\in\{1,2,\ldots,n\}$.
Proof.
The proof can be found in Appendix A.2. ∎
Since it is well known that the Hadamard matrix is highly effective at eliminating outliers Ashkboos et al. (2024b); Chee et al. (2024); Tseng et al. (2024), rotation is increasing the non-uniformity of the weight distributions. As presented in Fig. 3, we studied an empirical relation between the increase in the quantization error of output activation and the excess kurtosis of weight, after applying rotation. Clearly, the quantization error is generally enlarged when the excess kurtosis is increased by rotation. See Appendix A.3 for details and further discussion.
@@ 4 Methodology
(a) RCP first applies a Hadamard-based rotation to the weight and determines initial clipping values via grid search. The weight is then partitioned into 4 bins with learnable parameters.
(b) A diagram of Learnable Direct Partitioning.
@@
Figure 4: Illustration of Learnable Direct Partitioning (LDP) with rotation-aware clipping.
Our proposition is a QAT algorithm named Rotate, Clip, and Partition (RCP) that combines the idea of random rotation with our novel Learnable Direct Partitioning (LDP) quantizer.
Overall, RCP is a self-knowledge distillation Hinton et al. (2015) algorithm that solves the following optimization problem:
$\begin{split}\operatorname*{minimize}_{\Theta_{S}}\mathbb{E}_{(x,y)\sim\mathbb{D}}\left[\mathcal{D}_{CAKLD}(P_{\Theta_{T}}||P_{\Theta_{S}})\right],\\ \mathcal{D}_{CAKLD}(P_{\Theta_{T}}||P_{\Theta_{S}})=\alpha D_{KL}(P_{\Theta_{S}}||P_{\Theta_{T}})\\ +(1-\alpha)D_{KL}(P_{\Theta_{T}}||P_{\Theta_{S}}),\end{split}$ (3)
where $\Theta_{T}$ is the frozen full-precision teacher model, $\Theta_{S}$ is the student model quantized with LDP, $P_{\Theta}$ is the logit distribution produced by a model $\Theta$, $\mathbb{D}$ is the training dataset containing pairs of input text $x$ and label text $y$. The $\mathcal{D}_{CAKLD}$ is the confidence-aware KL divergence loss adopted from BitDistiller Du et al. (2024) with the empirical confidence $\alpha$ measured on $\Theta_{T}$.
4.1 Rotate: Applying Hadamard Transforms
The first step of our method is to apply (randomized) Hadamard transforms to model weights, following rotation-based PTQ algorithms Ashkboos et al. (2024b); Liu et al. (2024b); Lin et al. (2024a). We formulate this procedure as follows:
$\mathrm{\mathbf{W}}_{r}=\mathbf{R}^{T}_{front}\mathrm{\mathbf{W}}\mathbf{R}_{rear},$ (4)
where $\mathbf{R}^{T}_{front}$ and $\mathbf{R}_{rear}$ are Hadamard matrices multiplied to the front and rear side of a model weight $\mathrm{\mathbf{W}}$, respectively. The choice of $\mathbf{R}^{T}_{front}$ and $\mathbf{R}_{rear}$ can be identified in Fig. 2.
Note that the $\mathrm{\mathbf{W}}_{r}$ is pre-computed before any optimization to increase memory efficiency and better QAT performance. For further details and ablation, see Appendix A.5.
4.2 Clip: Learnable Clipping with Grid-Search Initialization
Clipping is an essential technique to limit quantization range via clamping out extreme values from the maximum and minimum sides of the model weights Lin et al. (2024b); Shao et al. (2024); Choi et al. (2018); Esser et al. (2020).
OmniQuant Shao et al. (2024) introduces learnable weight clipping (LWC) to dynamically determine the optimal quantization range by modifying the static quantization range $h$ in Eqn. 1 as follows:
$h_{lwc}=\sigma(\gamma)\mathrm{max}(\mathrm{\mathbf{W}_{r}})-\sigma(\beta)\mathrm{min}(\mathrm{\mathbf{W}_{r}}),$ (5)
where $\beta$ and $\gamma$ are learnable parameters allocated for each quantization unit and $\sigma$ is the sigmoid function.
To enhance stability of RCP, we find the initial point of the clipping parameters $\beta^{*},\gamma^{*}$ in a rotation-aware manner, based on the grid-search strategy Lin et al. (2024b) that solves the following problem on a small calibration dataset $\mathcal{D}_{cal}$:
$\begin{split}\beta^{*},\gamma^{*}=\operatorname{argmin}_{\beta,\gamma}\|Q(\mathrm{\mathbf{W}}_{rc})\textbf{X}_{r}-\mathrm{\mathbf{W}}_{rc}\textbf{X}_{r}\|^{2},\\ \mathrm{\mathbf{W}}_{rc}=\mathrm{clip}(\mathrm{\mathbf{W}}_{r},\sigma(\beta)\mathrm{min}(\mathrm{\mathbf{W}_{r}}),\sigma(\gamma)\mathrm{max}(\mathrm{\mathbf{W}_{r}})),\end{split}$ (6)
where $Q$ is the quantization function defined in Eqn. 1, $\mathrm{\mathbf{W}}_{rc}$ is the rotated and clipped weight, and $\mathrm{\mathbf{X}}_{R}$ is the rotated activation. In subsequent procedures, $\beta$ and $\gamma$ are learned via backpropagation, constantly searching for optimal dynamic quantization range on given data and updated model weights.
4.3 Partition: Learnable Direct Partitioning
The main goal of this work is to design a method to realize a non-uniform integer quantizer that learns from data. Prior arts such as N2UQ Liu et al. (2022) and LLT Wang et al. (2022) combine non-uniform quantization and uniform dequantization schemes to benefit from increased representational capability while being hardware friendly, however, we find this scheme results in suboptimal performance. Instead, LDP performs both steps in a non-uniform fashion to minimize the impact of errors from extreme weight quantization.
Non-uniform quantization via partitioning
The core idea of LDP is to partition the dynamic quantization range in a differentiable way by introducing two learnable parameters $s_{1}$ and $s_{2}$ per quantization unit. By applying sigmoid function to them, LDP directly splits $h_{lwc}$ into three partitions:
$\begin{split}p_{1}=\sigma(s_{1}),\\ p_{2}=(1-p_{1})\sigma(s_{2}),\\ p_{3}=(1-p_{1})(1-\sigma(s_{2})),\end{split}$ (7)
where $s_{1}$ takes the left $\sigma(s_{1})*100$% of $h_{lwc}$, $s_{2}$ takes the left $\sigma(s_{2})*100$% of the remaining range $(1-p_{1})h_{lwc}$, and the last partition is determined trivially.
We set the quantization grid $\{t_{i}|i\in\{1,2,\dots,2^{N}-1\}\}$ at the center of each partition as follows:
$t_{i}=t_{i-1}+\frac{p_{i-1}+p_{i}}{2},\;\mathrm{where}\;t_{1}=p_{1}/2.$ (8)
This obtains the quantization function of LDP as follows:
$Q_{LDP}(\textbf{W}_{rc})=\text{u}(\frac{\textbf{W}_{rc}}{h_{lwc}}-t_{1})+\text{u}(\frac{\textbf{W}_{rc}}{h_{lwc}}-t_{2})+\text{u}(\frac{\textbf{W}_{rc}}{h_{lwc}}-t_{3}),$ (9)
where $u(x)$ is the step function.
We rationalize this formulation in three points: 1) the dynamic range $h_{lwc}$ is seamlessly filled out since $\sum_{i=1}^{3}p_{i}=1$ is guaranteed; 2) each partition is constrained between 0% and 100% as the sigmoid re-parametrization ensures each $p_{i}\in[0,1]$; 3) no matter how the partitioning parameters are updated, the ordering of the partitions stays the same by the construction of $p_{i}$.
The initialization of $s_{1}$ and $s_{2}$ is set to $\sigma^{-1}(1/3)$ and $\sigma^{-1}(1/2)$, respectively (i.e., the dynamic range is uniformly partitioned at the beginning). Technically, the grid-search strategy Lin et al. (2024b) can also be employed to jointly find the optimal partitioning parameters; however, the computational cost will grow exponentially since we have to iterate over a four-dimensional optimization loop (two for LWC and two for LDP).
Non-uniform dequantization
The overall design of the quantization function in Eqn. 9 is imported, and the dequantization function of LDP is given by:
$\begin{split}DQ_{LDP}(\mathrm{\mathbf{W}}_{rc})=&\sigma(\beta)\text{min}(\textbf{W}_{rc})\\ +h_{lwc}\biggl{(}&\text{u}(\frac{\textbf{W}_{rc}}{h_{lwc}}-t_{1})(w_{1}-w_{0})\\ +&\text{u}(\frac{\textbf{W}_{rc}}{h_{lwc}}-t_{2})(w_{2}-w_{1})\\ +&\text{u}(\frac{\textbf{W}_{rc}}{h_{lwc}}-t_{3})(w_{3}-w_{2})\biggr{)},\end{split}$ (10)
where the dequantization grid $\{w_{i}|i\in\{0,1,2,3\}\}$ is additionally introduced in Eqn. 11.
$w_{i}=\begin{cases}0,&\text{if }i=0\\ \frac{t_{i}+t_{i+1}}{2},&\text{if }0<i<3\\ 1,&\text{if }i=3.\end{cases}$ (11)
This means that the full-precision weight elements whose normalized value is smaller than the first quantization grid (i.e., $\mathrm{W}/h_{lwc}<t_{1}$) are mapped to the minimum possible value in the dynamic range $\sigma(\beta)\mathrm{min}(\mathrm{\mathbf{W}_{rc}})$. Likewise, the elements that satisfy $\mathrm{W}/h_{lwc}>t_{3}$ are mapped to the maximum value $\sigma(\gamma)\mathrm{max}(\mathrm{\mathbf{W}_{rc}})$ [^3] Since $\sigma(\beta)\mathrm{min}(\mathrm{\mathbf{W}_{rc}})+h_{lwc}=\sigma(\gamma)\mathrm{max}(\mathrm{\mathbf{W}_{rc}})$. The others in the middle are mapped to the center of the second and third quantization bin, realizing non-uniform dequantization.